OpenStudy (anonymous):

how to find the fundamental period of 15cos(4x) + 8

OpenStudy (dumbcow):

period = 2pi/4 = pi/2

OpenStudy (anonymous):

divide the period of cosine by the coefficient of x and you're done. There are many properties of periodic functions. It'd be better if you do a little study on that ,always helpful.

OpenStudy (ybarrap):

Angular frequency is \(\tt \omega=\dfrac{2\pi}{T} \), where T is the period. So \(\tt T= \dfrac{2\pi}{\omega}\). T is the fundamental frequency. You have \( A\cos(\omega t) + 8\), with \( A=15\) and \(\omega=4\)

OpenStudy (ybarrap):

T is the period, not the frequence. f = 1/T is the fundamental frequency.